if 8x+5x+2x+4x=114
then 5x+3=?
please help solve

Answer by stanbon(75887)

if 8x+5x+2x+4x = 114
19x = 114
x = 114/19
x = 6
----
then 5x+3=?
5*6+3 = 33
=========================
Cheers,
Stan H.